OpenStudy (anonymous):

Change 16/50 to a percent

Directrix (directrix):

percent means hundredths. 16/ 50 * 2/2 = 32 / 100 = 32 %

OpenStudy (anonymous):

Do you have to do this all by hand? Do long division and divide 16 by 50. You will get a decimal. What do you do to convert a decimal to a percent? Move the decimal point over 2 places to the right (ALWAYS).

OpenStudy (anonymous):

On the calculator, 16 divided by 50 is 0.32 That is a decimal number. So move the decimal point over 2 places to the right and you will get 32%.

Directrix (directrix):

I didn't do the 16/ 50 by hand. Because percent means hundredths, I doubled the 50 in the denominator and the numerator. 32/100 means 32 percent. Yes. What do you do to convert a decimal to a percent? Move the decimal point over 2 places to the right (ALWAYS). Question: What would 5 be as a percent?

OpenStudy (anonymous):

It would be 500%

OpenStudy (anonymous):

if 5 is a decimal number

Directrix (directrix):

Yes. Interesting to me is that percentage is commutative (order doesn't matter). 10% of 100 = 100% of 10. Cool.

OpenStudy (anonymous):

Let's say something was at 200 and went up to 1000. 1000/200 equals 5. That's considered a decimal number in this case and so that means it went up 500%

OpenStudy (anonymous):

10% of 100 is 10.

OpenStudy (anonymous):

so if something is at 200 and it increases to 400, it went up 100%. If something is at 200 and it goes up to 600, it increased by 300%.

Directrix (directrix):

Increases seems like an additive operation. It is tricky. I interpret these words as "false." at 200 and it goes up to 600, it increased by 300%. because 200 + 300% of 200 = 200 + 3(200) = 200 + 600 = 800 and not 600.

OpenStudy (anonymous):

Yeah I guess it can be tricky. Because if it went up 25%, you'd say 200*1.25 (the one accounts for the 200 and the 0.25 counts for the 25% increase). So it would be 250 if 200 went up by 25%.

OpenStudy (anonymous):

so if it stays at 200, then it went up by 1 (the decimal number). 200*1 = 200

Directrix (directrix):

The same sort of tricky business comes when saying decreased by 50% or some other. If it stays at 200, it increased by 0 which is 0%.

OpenStudy (anonymous):

Staying the same means multiplying it by 1. Any other number you multiply it by means either an increase or decrease in percentage. Try it out as an exercise.

Directrix (directrix):

But, adding 1 is not staying the same.

OpenStudy (anonymous):

Actually, I told you the wrong thing for these percentage increases and I feel so guilty for it. Okay, here's the method if you have an increase problem. Say 200 increases to 300. Here's what you do: 1. Subtract the initial number you started out with from the final number. 300 - 200 = 100. 2. Divide that number you got by your initial number. 100/200 = 0.5 3. That last number you got was a decimal. Convert this into a percent. 0.5*100 = 50% increase. Again, I am SO SORRY for steering you in the wrong direction. But multiplying something by 1 means that it didn't go up or down. Multiplying it by any other number means that there was either an increase or decrease in percent.